A 100W, 250V lamp is connected in parallel with an unknown resistance ‘R,’ across a 250 V supply. The total power dissipated in the circuit is 1100V. Find the value of unknown resistance. Assume the resistance of the lamp remains unaltered. 


1
Expert's answer
2021-02-25T11:23:38-0500

power dissipated by the resistor"=1100W-100W=1000W"

"1000=\frac{v^{2}}{R}=\frac{250^{2}}{R}"

"R=\frac{62500}{1000}=62.5\Omega"